OpenStudy (anonymous):

Is this right? A salt truck drops 39 kilograms of salt per minute. How many grams of salt does the truck drop per second. 39kilograms/1minute (times) 1000milligrams/1gram (times) 1minute/60seconds=39000grams/60seconds=650grams/1second

OpenStudy (jhannybean):

So you already worked out the problem?

OpenStudy (jhannybean):

\[\sf \frac{39~kg}{1~min}~\times~\frac{10^3~g}{1~kg}~\times~ \frac{1~min}{60~sec} =\frac{39 \cdot 10^3~g}{60~sec} =\frac{39000~g}{60~sec} = \color{red}{\frac{650~g}{sec}} \] You're correct :)

OpenStudy (anonymous):

So I'm right? YAY! I had a hard time trying to figure it out. I just wanted to make sure if I did this right. Thank You! :D

OpenStudy (jhannybean):

No problem, good job \(\checkmark\)